# AI and the Net-Zero Journey: Energy Demand, Emissions, and the Potential for Transition

### v1

@@ 8 Mitigating AI’s Energy Impact: Strategies and Solutions
Research in green AI hardware can transform the sustainability of artificial intelligence by prioritizing energy efficiency and reducing environmental impact. Recent advancements include energy-efficient Graphics Processing Units (GPUs), Tensor Processing Units (TPUs), and neuromorphic chips, which consume significantly less power compared to traditional hardware [Clemm et al. (2024), Tabbakh et al. (2024)]. Advances in hardware design are now targeting the creation of AI accelerators that combine high performance with environmental sustainability. This effort includes developing processors that can run AI algorithms with very low energy consumption [Rahmani et al. (2023)]. Ongoing research aims to create AI devices capable of harvesting energy from their surroundings to power AI systems [Bolón-Canedo et al. (2024)]. Examples include utilizing ambient light, vibrations, and heat. By incorporating energy-harvesting features, these AI devices could become more self-sufficient, reducing reliance on external power sources and promoting a more sustainable deployment of AI [Divya et al. (2023), Ali et al. (2023)].
@@
Neuromorphic computing represents a significant shift in computing technology, inspired by the architecture of the human brain to create more efficient and adaptive systems. The human brain’s remarkable energy efficiency allows it to perform complex tasks with far less energy than conventional computers. By emulating this efficiency, neuromorphic computing aims to reduce the energy consumption of computational processes to levels similar to biological brains. This is achieved through the use of spiking neural networks (SNNs), which mimic the behavior of biological neurons, leading to substantial energy savings and performance enhancements, especially in AI applications that require real-time processing and decision-making. Additionally, a sustainability proposal by Oh et al. (2023) suggests creating biodegradable neural network hardware to mitigate the impact of technological waste.
@@
Powering data centers with renewable energy sources is a crucial step towards reducing the carbon footprint of the rapidly growing digital infrastructure. Integrating renewable energy, such as solar and wind power, into data center operations can help mitigate the environmental impact of these energy-intensive facilities [Brown and Gorham (2014)]. However, the variability of renewable energy sources poses challenges, which can be addressed through innovative solutions like virtual batteries that shift computational demand to match power availability [aga (2021)]. Additionally, advancements in energy storage technologies and smart grid integration are essential for ensuring a stable and reliable power supply [Hussain (2024)]. These efforts not only support sustainability goals but also align with the increasing corporate commitments to carbon neutrality [aga (2021)].
@@
CCS is a versatile technology and can be an enabler for data center decarbonization, with the potential to capture 90-95+ percent of a power generation facility’s CO2 emissions [Barlow et al. (2025)]. A business model is emerging which applies CCS to purpose-built power generation facilities "behind the meter" (not connected to municipal electric grids) to minimize data center carbon emissions while meeting their firm, high-reliability power demand. In this model, CO2 emissions from natural gas turbines are captured, transported, and permanently stored before they reach the atmosphere – significantly decarbonizing the data center power supply. Industry is responding to the data center decarbonization opportunity in the U.S. and other countries. Further details can be found in [Glo (2025)].
@@
Recent advancements in deep learning have focused on reducing the size of neural networks to enhance computational efficiency and enable deployment in resource-constrained environments. One approach involves a pruning algorithm that retrains the smaller, pruned model at its faster, initial learning rate, resulting in models that are simpler and more accurate [Renda et al. (2020)]. Another comprehensive review highlights techniques such as pruning, quantization, low-rank factorization, knowledge distillation, and transfer learning, all of which effectively reduce model size while retaining performance. A study on global weight compression demonstrates that deploying deep neural networks to devices requiring real-time processing can save memory, reduce storage size, and lower computational requirements [Dantas et al. (2024)]. An example advancement is the PocketNet paradigm, which aims to reduce the size of deep learning models used in medical image analysis. The authors propose a modification to existing convolutional neural network architectures that limits the growth of the number of channels, thereby significantly reducing the number of parameters. This approach maintains comparable performance to conventional neural networks while using up to 90 percent less GPU memory and speeding up training times by up to 40 percent. The paper demonstrates the effectiveness of PocketNet in various segmentation and classification tasks, highlighting its potential for deployment in resource-constrained settings [Celaya et al. (2023)].
@@
DeepSeek, a Chinese AI model, has garnered significant attention for its energy-efficient approach to artificial intelligence. Unlike traditional models, DeepSeek employs a "mixture of experts" technique during training, where only a portion of the model’s parameters are activated at any given time. This method, combined with improved reinforcement learning techniques, has led to more efficient training processes. However, while DeepSeek’s training phase is notably energy-efficient, its inference phase tends to be more energy-intensive. The model generates longer responses compared to similar models, which offsets some of the energy savings achieved during training [O’Donnell (2025)]. Despite its energy-efficient training, DeepSeek’s overall impact on energy consumption remains complex. The model’s longer inference responses require substantial computational power, which can lead to higher energy usage in data centers. This phenomenon is consistent with the broader trend in AI development, where efficiency gains in training often result in increased overall energy consumption due to the expanded capabilities and applications of the models [O’Donnell (2025)]. As companies strive to develop more intelligent systems, they tend to invest more in training, thereby using more energy. This dynamic underscores the importance of considering both training and inference phases when evaluating the energy efficiency of AI models. The introduction of DeepSeek has significant implications for the tech industry and energy sector. While its energy-efficient training methods offer a promising solution to the growing energy demands of AI, the model’s inference phase highlights the need for balanced approaches to AI development. The rapid adoption of DeepSeek-like models could strain power grids and challenge energy security.
@@ 9 Conclusions
We provided a probabilistic assessment of CO2 emissions based on the mean scenario derived from recent projections by the 2024 (2025); Paccou and Wijnhoven (2024) using the methodology presented in She (2025) . Our P5, P50, and P95 cases indicate that while AI growth is expected to continue over the next ten years, it is unlikely to reach extreme scenarios depicted in 2024 (2025); Paccou and Wijnhoven (2024). Instead, we anticipate deviations from the mean scenario, but not to such extreme extents. This is because most emissions are expected to come from data centers, and the growth in data centers does not support such extreme scenarios. Additionally, we do not foresee an energy crisis scenario, as there has been no significant precedent for such an event in the past 20 years, including during the COVID-19 pandemic.

### v2

@@ 8 Mitigating AI’s Energy Impact: Strategies and Solutions
Research in green AI hardware can transform the sustainability of artificial intelligence by prioritizing energy efficiency and reducing environmental impact. Recent advancements include energy-efficient Graphics Processing Units (GPUs), Tensor Processing Units (TPUs), and neuromorphic chips, which consume significantly less power compared to traditional hardware [clemm2024greenaicurrentstatus, tabbakh2024]. Advances in hardware design are now targeting the creation of AI accelerators that combine high performance with environmental sustainability. This effort includes developing processors that can run AI algorithms with very low energy consumption [Rahmani2023]. Ongoing research aims to create AI devices capable of harvesting energy from their surroundings to power AI systems [BOLONCANEDO2024128096]. Examples include utilizing ambient light, vibrations, and heat. By incorporating energy-harvesting features, these AI devices could become more self-sufficient, reducing reliance on external power sources and promoting a more sustainable deployment of AI [DIVYA2023108084, ALI2023101124].
@@
Neuromorphic computing represents a significant shift in computing technology, inspired by the architecture of the human brain to create more efficient and adaptive systems. The human brain’s remarkable energy efficiency allows it to perform complex tasks with far less energy than conventional computers. By emulating this efficiency, neuromorphic computing aims to reduce the energy consumption of computational processes to levels similar to biological brains. This is achieved through the use of spiking neural networks (SNNs), which mimic the behavior of biological neurons, leading to substantial energy savings and performance enhancements, especially in AI applications that require real-time processing and decision-making. Additionally, a sustainability proposal by Oh2023 suggests creating biodegradable neural network hardware to mitigate the impact of technological waste.
@@
Powering data centers with renewable energy sources is a crucial step towards reducing the carbon footprint of the rapidly growing digital infrastructure. Integrating renewable energy, such as solar and wind power, into data center operations can help mitigate the environmental impact of these energy-intensive facilities [brownandgorham2014]. However, the variability of renewable energy sources poses challenges, which can be addressed through innovative solutions like virtual batteries that shift computational demand to match power availability [agarwal2021]. Additionally, advancements in energy storage technologies and smart grid integration are essential for ensuring a stable and reliable power supply [hussain2024]. These efforts not only support sustainability goals but also align with the increasing corporate commitments to carbon neutrality [agarwal2021].
@@
CCS is a versatile technology and can be an enabler for data center decarbonization, with the potential to capture 90-95+ percent of a power generation facility’s CO2 emissions [Barlow2025]. A business model is emerging which applies CCS to purpose-built power generation facilities "behind the meter" (not connected to municipal electric grids) to minimize data center carbon emissions while meeting their firm, high-reliability power demand. In this model, CO2 emissions from natural gas turbines are captured, transported, and permanently stored before they reach the atmosphere – significantly decarbonizing the data center power supply. Industry is responding to the data center decarbonization opportunity in the U.S. and other countries. Further details can be found in [GlobalCCSInstitute2025].
@@
Recent advancements in deep learning have focused on reducing the size of neural networks to enhance computational efficiency and enable deployment in resource-constrained environments. One approach involves a pruning algorithm that retrains the smaller, pruned model at its faster, initial learning rate, resulting in models that are simpler and more accurate [renda_2020]. Another comprehensive review highlights techniques such as pruning, quantization, low-rank factorization, knowledge distillation, and transfer learning, all of which effectively reduce model size while retaining performance. A study on global weight compression demonstrates that deploying deep neural networks to devices requiring real-time processing can save memory, reduce storage size, and lower computational requirements [dantas_2024]. An example advancement is the PocketNet paradigm, which aims to reduce the size of deep learning models used in medical image analysis. The authors propose a modification to existing convolutional neural network architectures that limits the growth of the number of channels, thereby significantly reducing the number of parameters. This approach maintains comparable performance to conventional neural networks while using up to 90 percent less GPU memory and speeding up training times by up to 40 percent. The paper demonstrates the effectiveness of PocketNet in various segmentation and classification tasks, highlighting its potential for deployment in resource-constrained settings [celaya_2024].
@@
DeepSeek, a Chinese AI model, has garnered significant attention for its energy-efficient approach to artificial intelligence. Unlike traditional models, DeepSeek employs a "mixture of experts" technique during training, where only a portion of the model’s parameters are activated at any given time. This method, combined with improved reinforcement learning techniques, has led to more efficient training processes. However, while DeepSeek’s training phase is notably energy-efficient, its inference phase tends to be more energy-intensive. The model generates longer responses compared to similar models, which offsets some of the energy savings achieved during training [odonnell2025]. Despite its energy-efficient training, DeepSeek’s overall impact on energy consumption remains complex. The model’s longer inference responses require substantial computational power, which can lead to higher energy usage in data centers. This phenomenon is consistent with the broader trend in AI development, where efficiency gains in training often result in increased overall energy consumption due to the expanded capabilities and applications of the models [odonnell2025]. As companies strive to develop more intelligent systems, they tend to invest more in training, thereby using more energy. This dynamic underscores the importance of considering both training and inference phases when evaluating the energy efficiency of AI models. The introduction of DeepSeek has significant implications for the tech industry and energy sector. While its energy-efficient training methods offer a promising solution to the growing energy demands of AI, the model’s inference phase highlights the need for balanced approaches to AI development. The rapid adoption of DeepSeek-like models could strain power grids and challenge energy security.
@@ 9 Conclusions
We provided a probabilistic assessment of CO2 emissions based on the mean scenario derived from recent projections by the iea2025; paccouandwijnhoven2024 using the methodology presented in Shell_Sce_2025 . Our P5, P50, and P95 cases indicate that while AI growth is expected to continue over the next ten years, it is unlikely to reach extreme scenarios depicted in iea2025; paccouandwijnhoven2024. Instead, we anticipate deviations from the mean scenario, but not to such extreme extents. This is because most emissions are expected to come from data centers, and the growth in data centers does not support such extreme scenarios. Additionally, we do not foresee an energy crisis scenario, as there has been no significant precedent for such an event in the past 20 years, including during the COVID-19 pandemic.
@@
10 Acknowledgements
The authors thank Shell Information Technology International Inc. and Shell International Exploration and Production Inc. for permission to publish this work. Additionally, we are grateful to our colleagues—including Juben Chheda, Mariela Araujo, Christian Davies, Pak Leung, and Albena Mateeva—for their meaningful discussions and valuable contributions that shaped the final paper.
Table 1: The anticipated electricity consumption for various tasks related to AI chatbot models [Luccioni20240605]. As shown, text-based tasks are significantly less demanding in terms of electricity consumption, while image-based tasks require substantially more energy. For context, generating 1,000 words of text is comparable to charging a smartphone battery by 9%. In contrast, completing image-based tasks can consume as much energy as charging approximately 522 smartphones.
Task Inference energy consumption (Wh)
Text classification 2
Text generation 47
Summarization 49
Object detection 38
Image captioning 63
Image generation 2907
Table 2: The list of key variables potentially impacting the CO2 emissions at data centers include the following factors: electricity consumption by the semiconductor industry [greenpeacereport2024], electricity consumption by data centers [paccouandwijnhoven2024], the energy mix factor indicating the percentage of fossil feedstock-based electricity generation, the percentage of AI workloads at the data centers, and CO2 emissions in Mt per year. The data from Figure 7b were used to find the mean scenario and the corresponding numbers.
Year
Electricity consumption by
semiconductor industry (TWh)
Electricity consumption by
data centers (TWh)
Energy mix factor
AI workload
at data centers
CO2 emissions
(Mt/year)
2020 91 269 0.62 0.02 1.03
2021 101 300 0.59 0.03 2.07
2022 112 330 0.59 0.06 4.12
2023 126 361 0.56 0.11 8.18
2024 140 416 0.55 0.17 14.18
2025 156 482.5 0.53 0.2 20.08
2026 173 550 0.52 0.3 32.7
2027 191 650 0.50 0.43 54.7
2028 210 746 0.47 0.55 80
2029 229 838 0.45 0.65 104.5
2030 240 918 0.43 0.73 126
2031 245 981 0.41 0.69 124
2032 266 1038 0.39 0.67 124
2033 288 1080 0.37 0.65 124
2034 310 1118 0.35 0.63 121
2035 334 1148 0.33 0.64 123